OpenStudy (anonymous):

find an equation of line that passes through through the point 2,4 and has a slope of -1

OpenStudy (anonymous):

y=ax+b a=slope=-1 y=-x+b 4=-1*2+b 4=-2+b b=6 y=-x+6